OpenStudy (anonymous):

Can anyone help me with finding the domain and range for this quadratic equation y = –x^2 – 14x – 52?

OpenStudy (anonymous):

\[b^2 -4ac \geq 0\]

OpenStudy (anonymous):

What..?

OpenStudy (anonymous):

Take it to the completed square form

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

Factorize it

OpenStudy (anonymous):

y = –x^2 – 14x – 52 = -(x^2+14x)-52 =-(x^2+2*x*7+7^2-7^2)-52 = -{(x+7)^2-49}-52 =-(x+7)^2-3

OpenStudy (anonymous):

thats the completed square form

OpenStudy (anonymous):

Now, domain is x E R.......

OpenStudy (anonymous):

For range, (x+7)^2>=0 -(x+7)^2<=0 -(x+7)^2-3<=-3 y<=-3

OpenStudy (anonymous):

got it?

OpenStudy (precal):

I always look at the graphs when determining the domain and range

OpenStudy (precal):

https://www.desmos.com/calculator

OpenStudy (precal):

domain is all real numbers range is all numbers less than and equal to -3 just like sauravshakya stated

OpenStudy (precal):

use the above link if you do not own a graphing calculator

OpenStudy (anonymous):

Domain is R becoz...there is no...Restrictions...)) in the given equation

OpenStudy (precal):

yes and domain is all real numbers because all of the x's are being used on the number line to plot the function